# Governing Dynamic Capabilities: Cryptographic Binding and Reproducibility Verification for AI Agent Tool Use

### 15 Mar 2026 v1

@@ Abstract
[…]
@@ 1.1 Agent Governance: A Missing Foundation
AI agents are no longer isolated tools. Through protocols like MCP (Model Context Protocol) [8] and A2A (Agent-to-Agent Protocol) [8], modern agents dynamically discover tools, invoke external APIs, execute code, and orchestrate other agents at runtime. The MCP ecosystem alone has grown to over 1,000 tool servers in under a year [32]. As these agents assume real-world responsibilities—managing enterprise credentials, executing financial transactions, accessing sensitive data—a fundamental question emerges: what properties must an agent ecosystem satisfy for humans to trust it?
This question has been raised from multiple directions. The EU AI Act (Regulation 2024/1689) mandates traceability, logging, and conformity assessment for high-risk AI systems (Articles 12, 43, 72). NIST’s AI Agent Standards Initiative [26] identifies security controls, governance oversight, and monitoring as key requirements. The TRiSM framework [28] structures agent governance around explainability, security, privacy, and lifecycle management. Schroeder de Witt [11] calls for secure interaction protocols that prevent insecure emergent behaviors. Cheng et al. [43] ground agent safety in transparency, accountability, and trustworthiness. Allegrini et al. [44] formalize 31 safety and liveness properties for agentic task lifecycles in temporal logic. Ghosh et al. [45] propose an operational safety framework with red-teaming methodology, policy-based tool scoping, and optional trace signing—but without cryptographic capability binding, formal adversary models, or verifiable replay mechanisms.
Yet despite this convergence of concern, no existing framework translates these governance aspirations into enforceable cryptographic infrastructure. The gap is not in recognizing what is needed, but in building it. We propose three Agent Governance Requirements that distill the common thread across these efforts into technically precise, verifiable properties:
@@
G2: Behavioral Verifiability. An agent’s actions within its authorized capability envelope can be independently verified post-hoc, enabling detection of model substitution, unauthorized tool use, or output manipulation.
@@
These requirements are not novel in aspiration—they reflect the consensus of legal [33, 34], regulatory [26], and technical [11, 28, 43, 44, 45] communities. What is novel is the claim that all three must hold simultaneously, and that none is achievable with existing agent infrastructure. G1 alone (without G2 and G3) permits certified agents to misbehave undetectably within their capability envelope. G2 alone (without G1) cannot distinguish authorized from unauthorized capabilities. G3 alone (without G1 and G2) produces audit logs of unverified actions by unverified agents.
1.2 Why the Current Landscape Falls Short
The current agent security landscape can be mapped into four layers (Figure 1), each addressing fragments of the governance requirements but none satisfying them:
Identity platforms (Keyfactor [19], SPIFFE/SPIRE [21, 22]) issue certificates binding an agent to a public key—but a certificate remains valid regardless of how many tools the agent acquires after issuance. G1 violated: capability changes are invisible to the identity layer.
Authorization frameworks (OIDC-A [4], Authenticated Delegation [3], Cerbos [39]) define what an authenticated agent may access—but authorization is evaluated against declared capabilities at grant time, with no mechanism to detect drift. G1 violated: authorization does not track capability changes.
Runtime gateways (Gravitee [40], Lasso Security [41]) inspect tool calls in real time—but are fundamentally stateless, with no memory of prior tool configurations and no behavioral verification. G2 and G3 violated.
Governance platforms (Credo AI [42]) maintain agent registries documenting capabilities—but operate at the documentation layer without cryptographic enforcement. G1 violated: declared capabilities are not bound to the agent’s actual configuration.
The root cause is a failure to recognize a fundamental architectural distinction in LLM-based agents. Inside a transformer’s forward pass, tool definitions and user context are indistinguishable token sequences. But at the orchestration layer, tool definitions determine which real-world actions the model’s outputs can trigger, while runtime context determines which actions the model chooses. This yields a principled capability-context separation: the capability envelope (model identity, tool definitions, permission scopes) governs what is possible and changes infrequently; the runtime context (conversation history, user instructions, memories) governs what is chosen and changes per-interaction. The security implication is direct: capability changes require re-authorization (G1); context changes require verifiability (G2) and auditability (G3).
Existing frameworks conflate the two. A tool addition and a new user message are treated identically—neither triggers re-evaluation. We call this the capability-identity gap: existing frameworks authenticate who an agent is and authorize what it may access, but none detect when the agent’s capabilities have changed. This gap enables a class of attacks we term silent capability escalation.
1.3 Our Framework: Bridging the Gap
The capability-context separation provides the theoretical foundation for three mechanisms, each targeting a specific governance requirement:
@@ 1.3 Our Framework: Bridging the Gap
Together, these three mechanisms provide what the existing landscape lacks: a stateful, cryptographic governance layer satisfying G1–G3, bridging identity platforms and runtime gateways. Our contribution is not in novel cryptographic primitives; rather, we show that standard mechanisms—X.509 certificates, SHA-256 hashing, Ed25519 signatures, hash chains—when composed through the capability-context separation, are sufficient to enforce G1–G3 with negligible overhead, providing actionable design guidance for how agents should be built, how they should interact, and how their behavior should be audited.
@@ 1.4 Threat Model and Security Guarantees
We consider a PPT adversary who can compromise up to $k$ agent nodes, create malicious agents, inject or modify tools, and manipulate network traffic (Dolev-Yao model). Critically, the adversary can perform silent capability escalation. The adversary cannot compromise non-agent nodes or break standard cryptographic primitives.
Under this model, we define and prove nine security properties:
  1. 1.
    Capability Binding: tool configuration change $\Rightarrow$ access denied (Theorem 5).
  2. 2.
    Trust Containment: child constraints $\leq_{\kappa}$ parent constraints (Theorem 4).
  3. 3.
    Chain Integrity: valid chain to trusted root required (Theorem 6).
  4. 4.
    Delegation Depth: chains bounded by ancestor’s $\mathit{max\_depth}$ (Theorem 7).
  5. 5.
    Reproducibility Soundness: different model/tools detected with overwhelming probability (Theorem 8).
  6. 6.
    Bounded Divergence: adversary-controlled model passing $n$ replay checks is $(n,\epsilon)$-indistinguishable from declared model (Theorem 9).
  7. 7.
    Chain Verifiability: a single unverifiable interior agent breaks end-to-end behavioral verification for all downstream nodes (Theorem 12).
  8. 8.
    Credential Isolation: colluding agents gain no extra permissions (Theorem 10).
  9. 9.
    Ledger Integrity: undetectable modification impossible (Theorem 11).
Properties 1–4 enforce G1. Properties 5–7 enforce G2 (with 7 characterizing its chain-level semantics). Properties 8–9 enforce G3.
@@ 1.5 Contributions
In summary:
  • We define three Agent Governance Requirements—capability integrity, behavioral verifiability, and interaction auditability—distilling converging regulatory, legal, and technical concerns into formally precise properties (§1.1).
  • We identify the capability-identity gap and the capability-context separation principle as its root cause and resolution (§1.2).
  • We show that standard cryptographic mechanisms—composed through the capability-context separation—are sufficient to enforce all three requirements with negligible overhead, providing design guidance for secure agent ecosystems.
  • We propose capability-bound agent certificates with monotonic trust propagation (§3).
  • We introduce reproducibility as a trust primitive and establish that reproducibility is a necessary condition for behavioral verifiability: high-assurance agent deployments must be designed for deterministic output to be auditable (§4).
  • We define a verifiable interaction ledger for tamper-evident forensic reconstruction (§5).
  • We prove that sufficient replay verification yields bounded divergence: an adversary-controlled model that passes $n$ replay checks is provably $(n,\epsilon)$-indistinguishable from the declared model, with $\epsilon$ exponentially small in nn4.7).
  • We establish chain verifiability: behavioral verifiability is a chain property, not a node property—a single unverifiable interior agent breaks end-to-end verification for all downstream nodes, yielding principled deployment constraints on model placement in trust trees (§4.8).
  • We implement a working MCP governance proxy that enforces all three requirements in real-time with $\sim$}120 µs overhead per call, detecting silent capability escalation, tool trojanization, and certificate forgery (§7).
  • We formalize and prove nine security properties under a realistic adversary model (§6).
  • We implement a Rust prototype achieving 97 µs verification, ${\sim}0.62$ ms governance overhead, and ${\sim}1{,}200{,}000\times$ speedup over BAID (§7).
  • We deploy the full governance stack in an end-to-end multi-agent pipeline (5–20 agents with real LLM calls), demonstrating 0.12% overhead, sub-linear per-agent scaling, and detection of all in-situ attacks—while OAuth 2.1 and OpenTelemetry baselines detect none (§7.4).
1.6 Paper Organization
§2 presents the threat model and attack scenarios. §3 defines capability-bound certificates and the trust propagation tree. §4 introduces reproducibility commitments. §5 describes the verifiable interaction ledger. §6 presents formal security properties and proofs. §7 describes implementation and evaluation. §8 discusses limitations. §9 covers related work. §10 concludes.
@@ 2.3 Motivating Attack Scenarios
ATK-2: Model Substitution. $\mathcal{A}$ replaces the certified model with a fine-tuned variant. Detected via model binding mismatch and replay verification divergence.
@@
Delegation Attacks. ATK-4: Phantom Delegation. Compromised agent (T2) issues T1 certificate to sub-agent. Prevented by monotonic trust propagation. ATK-5: Chain Forgery. Fabricated certificate chain. Prevented by signature verification at every level. ATK-6: Depth Overflow. Deep delegation hierarchy. Prevented by strictly decreasing depth counters. ATK-7: Credential Collusion. Two or more compromised agents at lower tiers (e.g., T3++T3) attempt to combine their credentials to access resources restricted to a higher tier (T1). Prevented by independent per-agent credential evaluation—each agent’s tier and permissions are verified in isolation, so colluding agents gain no additional permissions beyond their individual certificates (Theorem 10).
@@ 3.5 Integration with A2A Agent Cards
Agent Certificates are designed as non-invasive extensions of the A2A ecosystem via an optional extensions field in the Agent Card containing the certificate PEM, chain, and skills manifest hash.
4 Reproducibility Commitments
4.1 Reproducibility as a Necessary Condition
A key observation motivates our approach: behavioral verifiability (G2) is impossible without reproducibility. If an agent’s outputs cannot be independently regenerated—even approximately—then no post-hoc mechanism can distinguish authorized behavior from unauthorized behavior. This is not merely a limitation of our framework; it is a fundamental constraint on any verification scheme that does not rely on trusted execution environments.
This observation has a prescriptive implication for secure agent system design: high-assurance agent deployments must be engineered for reproducible output. Concretely, this means using deterministic sampling (temperature==0, fixed seeds), preferring structured outputs (code, JSON, SQL) over free-form prose where security-critical decisions are involved, and selecting inference providers that guarantee bitwise or near-bitwise reproducibility [6]. Agents that cannot commit to any reproducibility level receive a trust tier downgrade (§4.4), making the security cost of non-reproducibility explicit in the governance framework.

### 19 Mar 2026 v2

@@ Abstract
AI agents now dynamically acquire tools, orchestrate sub-agents, and transact across organizational boundaries—yet no existing security layer can verify what an agent is capable of, whether it executed what it claims, or what actually happened in a multi-agent interaction. We trace this governance vacuum to a missing architectural distinction: the capability-context separation. Inside a transformer’s forward pass, tool definitions and user context are indistinguishable token sequences; at the orchestration layer, they have fundamentally different security semantics—tool definitions determine which real-world actions are possible (and change infrequently), while runtime context determines which actions are chosen (and changes per interaction). Existing frameworks conflate the two, enabling silent capability escalation—agents acquiring tools without invalidating credentials—and leaving cross-agent interactions without verifiable provenance.
From this principle, we derive three Agent Governance Requirements: capability integrity (G1) governs the capability envelope, behavioral verifiability (G2) ensures agents executed their declared computational process, and interaction auditability (G3) provides tamper-evident records for the runtime context—together defining what an agent ecosystem must enforce, independent of how. We prove two structural results with prescriptive architectural consequences: the Chain Verifiability Theorem, establishing that behavioral verification is a chain property (one unverifiable interior agent breaks end-to-end verification for all downstream nodes), and the Bounded Divergence Theorem, which transforms replay-based verification into a probabilistic safety certificate ($\epsilon\leq 1-\alpha^{1/n}$). We validate the framework with two crypto-agnostic instantiations—basic (Ed25519, SHA-256, hash chains; 97 µs verify) and enhanced (BBS+ selective disclosure, Groth16 DV-SNARK; 13.8 ms verify)—both satisfying the same nine security properties. A multi-provider reproducibility study (9 models, 7 providers) reveals $5.8\times$ variance in inference determinism, directly connecting model characteristics to governance architecture. End-to-end evaluation over 5–20 agent pipelines with real LLM calls confirms <<0.02% governance overhead, real-time detection of 7 end-to-end attack scenarios (covering G1, G2, and G3) with zero false positives, and post-hoc forensic tracing of runtime behavioral attacks validated against the Nemotron-AIQ dataset.
@@ 1.1 Agent Governance: A Missing Foundation
AI agents are no longer isolated tools. Through protocols like MCP [8] and A2A [9], modern agents dynamically discover tools, invoke external APIs, and orchestrate other agents at runtime. As these agents assume real-world responsibilities—managing credentials, executing financial transactions, accessing sensitive data—a fundamental question emerges: what properties must an agent ecosystem satisfy for humans to trust it?
This question has been raised from multiple directions: regulatory mandates for traceability and conformity assessment (EU AI Act, Articles 12, 43, 72), standards initiatives identifying security controls and governance oversight [27, 29], and research calling for secure interaction protocols [12], transparency [44], formal safety properties [45], and operational safety frameworks [46]. Yet no existing framework translates these aspirations into enforceable cryptographic infrastructure. This is a fundamental architectural gap: current agents acquire tools dynamically, execute arbitrary models, and interact without verifiable records.
We propose three Agent Governance Requirements—the minimum conditions under which an agent ecosystem can be considered governed:
@@
G2: Behavioral Verifiability (Inference Authenticity). An agent’s declared computational process can be independently verified—ensuring that the agent actually executed what its credentials claim. This requirement is implementation-agnostic: it may be satisfied by hardware attestation, software-based replay, or zero-knowledge proofs.
@@
[…]
Table 1: Governance gap analysis: undefended attacks when one or two governance requirements are absent. ATK-9 (blame shifting) requires both G2 and G3 for full resolution. All 12 attacks are defined in §2.
Missing Undefended Attacks #
Single requirement absent
@@
Two requirements absent
[…]
A critical corollary: even hardware TEEs [53, 54, 55] do not eliminate the need for G1 and G3. A TEE-attested agent that silently acquires new tools remains a security threat (G1 violation); a TEE-attested pipeline without tamper-evident records remains unauditable (G3 violation). Our framework provides the governance layer that TEE, by itself, cannot.
1.2 Why the Current Landscape Falls Short
The current agent security landscape maps into four layers (Figure 1), none satisfying G1–G3: Identity platforms [20, 22, 23] bind agents to keys but certificates remain valid after capability changes (G1 violated). Authorization frameworks [4, 3, 40] evaluate permissions at grant time without detecting drift (G1 violated). Runtime gateways [41, 42] inspect tool calls statelessly (G2, G3 violated). Governance platforms [43] maintain registries without cryptographic enforcement (G1 violated).
The root cause is a failure to recognize a fundamental architectural distinction. Inside a transformer’s forward pass, tool definitions and user context are indistinguishable token sequences; but at the orchestration layer, tool definitions determine which real-world actions are possible, while runtime context determines which actions are chosen. This yields the capability-context separation: the capability envelope (model identity, tool definitions, permission scopes) changes infrequently and requires re-authorization (G1); the runtime context (conversation history, user instructions) changes per-interaction and requires verifiability (G2) and auditability (G3).
Existing frameworks conflate the two—a tool addition and a new user message are treated identically. On the capability side, this creates the capability-identity gap: no existing framework detects when an agent’s capabilities have changed, enabling silent capability escalation. On the context side, it means cross-agent interactions lack verifiable provenance—no framework provides tamper-evident records linking which agent did what, with which inputs, producing which outputs—the foundation for accountability in multi-agent workflows.
To our knowledge, the capability-context separation is the first architectural abstraction that explains why agent security frameworks systematically fail to address capability drift—and it is not merely descriptive but prescriptive: it dictates the governance boundaries any secure agent system must enforce. As a concrete illustration, BAID [2] achieves strong code-level authentication via zkVM, binding agent identity to a program commitment $C_{P}=\mathrm{CommitProg}(P)$. Yet $C_{P}$ does not encompass dynamically acquired tools: an agent’s program binary remains unchanged when new tools are discovered via MCP [8] or plugin registries, leaving the entire capability envelope outside BAID’s security perimeter. This is not an implementation oversight but a direct consequence of designing without the capability-context separation—the framework conflates code identity with capability identity. Our G1 requirement, derived from this abstraction, prescribes that governance must commit to the complete capability set (model + tools + permissions), not merely the executing code.
1.3 Our Framework: Bridging the Gap
The capability-context separation provides the foundation for three enforcement mechanisms that instantiate the governance contracts G1–G3:
[…]
Together, these mechanisms provide a stateful, cryptographic governance layer satisfying G1–G3.
Crypto-agnostic design. G1–G3 define what to enforce, not how. The framework is crypto-agnostic: each requirement specifies a functional contract with pluggable cryptographic primitives—analogous to TLS cipher suite negotiation. We present a basic instantiation (Ed25519, SHA-256, hash chains; ${\sim}97$ µs verification) and an enhanced instantiation (BBS+ selective disclosure for G1, DV-SNARK zero-knowledge verification for G2; 13.8 ms). Both satisfy G1–G3 with identical security semantics, differing only in privacy properties and cost. This separation of governance logic from cryptographic mechanism demonstrates that agent governance is a composable architecture where organizations select backends matching their threat model.
@@ 1.4 Threat Model and Security Guarantees
We consider a PPT adversary who can compromise up to $k$ agent nodes, create malicious agents, inject or modify tools, manipulate network traffic (Dolev-Yao model), and perform silent capability escalation; it cannot compromise non-agent nodes or break standard cryptographic primitives. Under this model, we establish seven security properties and two structural theorems: Capability Binding (Property 4), Trust Containment (Property 3), Chain Integrity (Property 5), and Delegation Depth (Property 6) enforce G1; Reproducibility Soundness (Property 7), Bounded Divergence (Theorem 3), and Chain Verifiability (Theorem 4) enforce G2; Credential Isolation (Property 8) and Ledger Integrity (Property 9) enforce G3.
@@ 1.5 Contributions
Our contribution is a new governance architecture—identifying what must be governed, where governance boundaries belong, and why a specific composition is both necessary and sufficient—not new cryptographic primitives. This follows the pattern of Certificate Transparency [56], which composed Merkle trees and gossip protocols (both well-known) into infrastructure that fundamentally changed web PKI operation. The insight was in identifying what to compose and where to deploy it, not in new primitives; CT was published at IEEE S&P precisely because the architectural contribution—public auditability of CA behavior—was independently valuable. We make an analogous claim: the capability-context separation, the three governance requirements, and the structural results that connect them are independently valuable regardless of the specific cryptographic instantiation. Concretely: 1. A new vulnerability class and architectural abstraction. We identify silent capability escalation—agents acquiring or modifying tools at runtime without triggering existing security mechanisms (§1.2)—and trace it to the capability-context separation: tool definitions and user context are indistinguishable inside a transformer but have fundamentally different security semantics at the orchestration layer [57, 58]. To our knowledge, this is the first architectural abstraction that identifies where governance boundaries must be placed in agent systems and why existing frameworks—including those with strong cryptographic foundations [2]—systematically miss capability drift.
2. Structural results with architectural consequences. The Chain Verifiability Theorem4.9) shows that a single unverifiable interior agent breaks end-to-end behavioral verification, prescribing high-determinism models at interior trust-tree positions. The Bounded Divergence Theorem4.8) establishes that an adversary passing $n$ replay checks is $(n,\epsilon)$-indistinguishable from the declared model ($\epsilon\leq 1-\alpha^{1/n}$), framing software G2 as a probabilistic approximation of TEE. The Agent Governance Trilemma8.2) proves that capability, performance, and security cannot be simultaneously maximized for Turing-complete agents (via Rice’s theorem).
3. Empirical discovery connecting model determinism to chain security. A reproducibility study (9 models, 7 providers, 15,120 comparisons) reveals $5.8\times$ variance in inference determinism, which—combined with Chain Verifiability—determines which trust-tree positions a model qualifies for.
4. Two crypto-agnostic instantiations. Basic (Ed25519, SHA-256; 97 µs) and enhanced (BBS+, DV-SNARK; 13.8 ms) instantiations both satisfy nine security properties (§6). End-to-end evaluation over 5–20 agent pipelines with real LLM calls confirms <<0.02% governance overhead and real-time detection of 7 E2E attack scenarios (covering G1, G2, and G3) with zero false positives (§7.5). An additional 9 structural governance attacks are validated as automated unit tests, and 3 runtime behavioral attacks are validated via post-hoc forensic tracing against Nemotron-AIQ traces (§7).
@@ 2.3 Motivating Attack Scenarios
ATK-2: Model Substitution. $\mathcal{A}$ replaces the certified model with a fine-tuned variant. Detected via G2: hardware TEE attestation (deterministic), software replay verification divergence, or DV-SNARK CharMatch failure.
@@
Delegation Attacks. ATK-4: Phantom Delegation. Compromised agent (T2) issues T1 certificate to sub-agent. Prevented by monotonic trust propagation. ATK-5: Chain Forgery. Fabricated certificate chain. Prevented by signature verification at every level. ATK-6: Depth Overflow. Deep delegation hierarchy. Prevented by strictly decreasing depth counters. ATK-7: Credential Collusion. Two or more compromised agents at lower tiers (e.g., T3++T3) attempt to combine their credentials to access resources restricted to a higher tier (T1). Prevented by independent per-agent credential evaluation—each agent’s tier and permissions are verified in isolation, so colluding agents gain no additional permissions beyond their individual certificates (Property 8).
@@ 3.4 Certificate Verification Algorithm
3.5 Enhanced G1: Selective Disclosure via BBS+ Signatures
The basic G1 instantiation (Ed25519 + SHA-256) requires full disclosure: a verifier must see the entire skills manifest to recompute the hash and check capability binding. This suffices for single-organization deployments but is problematic in cross-organizational settings where revealing the full tool manifest leaks competitive intelligence (e.g., which proprietary tools an agent uses, its model version, or its rate limits).
BBS+ signatures [51] enable a strictly stronger property: an agent can prove that its certificate is valid and that specific capability predicates hold, without revealing the full certificate contents. BBS+ is a W3C Verifiable Credentials standard primitive with security reducing to the $q$-SDH assumption in bilinear groups [52].
Definition 5 (BBS+ Capability Certificate).
The issuer signs the certificate attributes as a vector of $L$ messages:
$\mathbf{m}=(m_{1},\ldots,m_{L})=(\mathit{id}_{v},\mathit{pk}_{v},\mu_{v},h_{1},\ldots,h_{n},\kappa_{v},\rho_{v},t_{s},t_{e})$
producing a BBS+ signature $\Sigma=(A,e,s)$ over $\mathbf{m}$.
Selective disclosure. To prove capability binding to a verifier, the agent derives a zero-knowledge proof of knowledge $\pi$ that:
  1. 1.
    The agent possesses a valid BBS+ signature $\Sigma$ from a trusted issuer.
  2. 2.
    The revealed subset $\mathbf{m}_{R}\subset\mathbf{m}$ (e.g., only the tool IDs, not model version or rate limits) matches the disclosed values.
  3. 3.
    The hidden attributes $\mathbf{m}_{H}=\mathbf{m}\setminus\mathbf{m}_{R}$ satisfy declared predicates (e.g., $\mathit{max\_tier}\geq 2$, $\mathit{model}\in\{\texttt{gpt-4o},\texttt{claude-4}\}$).
The verifier learns only the revealed attributes and the predicate outcomes—not the hidden values. This enables scenarios such as: an agent proves it has some valid tool manifest without revealing which tools; or proves its trust tier exceeds a threshold without revealing its exact tier or delegation depth.
Capability binding under BBS+. The capability binding invariant is preserved: at runtime, the agent recomputes $H_{v}(t)=\mathrm{SHA\text{-}256}(\mathrm{canonical}(S_{v}(t)))$ and includes a ZKP that $H_{v}(t)$ matches the committed hash in the signed message vector. Any tool change invalidates the proof.
Cost-privacy tradeoff. BBS+ selective disclosure proof generation costs ${\sim}14.7$ ms and verification ${\sim}13.8$ ms (benchmarked in §7), compared to ${\sim}97$ µs for basic Ed25519 verification—a ${\sim}142\times$ overhead. This is the concrete cost of privacy: the governance contract (G1) is identical; the privacy properties differ. Deployments choose between basic and enhanced G1 based on whether cross-organizational privacy justifies the latency cost. Crucially, these costs apply per-request only when selective disclosure is needed; for routine intra-organizational use, the basic instantiation suffices.
Security. BBS+ signature unforgeability reduces to the $q$-SDH assumption in Type-3 bilinear groups; the zero-knowledge property reduces to DDH. Both are standard assumptions with well-studied cryptanalytic history. The enhanced instantiation inherits all nine security properties of the basic instantiation, with the additional property of certificate privacy: no PPT verifier can extract hidden attributes from a selective disclosure proof.
@@ 3.6 Integration with A2A Agent Cards
Agent Certificates are designed as non-invasive extensions of the A2A ecosystem via an optional extensions field in the Agent Card containing the certificate PEM, chain, and skills manifest hash. The enhanced instantiation adds a disclosure_mode field (full or selective) and, for selective disclosure, the BBS+ proof $\pi$ replacing the raw certificate.
4 Inference Authenticity and Reproducibility
4.1 G2 as Inference Authenticity
G2 requires inference authenticity: verifying that an agent actually executed its declared model on the recorded inputs. This is an abstract governance requirement, independent of how it is achieved. Three implementation paths exist, each with distinct trust assumptions:
Hardware attestation (TEE). Trusted execution environments [53, 54] guarantee that the declared code was what actually executed. This provides the strongest G2 guarantee—deterministic, not probabilistic—but requires specialized hardware infrastructure (Intel SGX/TDX [53], ARM TrustZone [54], NVIDIA H100 Confidential Computing [55]) that is unavailable in most current LLM API deployments. TEE satisfies G2 but does not satisfy G1 or G3: a TEE-attested agent can still silently acquire tools (G1 violation), and TEE provides per-execution integrity without tamper-evident cross-agent audit trails (G3 violation). Our framework provides the governance layer that TEE alone cannot.
Software replay verification. When TEE is unavailable—the dominant case for cloud-hosted LLM APIs—G2 can be approximated through reproducibility-based replay: re-execute the agent’s declared model on recorded inputs and compare outputs. This provides a probabilistic software approximation of TEE, with guarantees formalized in Property 1 and Theorem 1. The approximation quality improves with the verification budget $n$.
Zero-knowledge verification (DV-SNARK, §4.10). A privacy-preserving variant of replay verification: the prover demonstrates output consistency without revealing prompts or outputs. This adds privacy guarantees to the software path at the cost of computational overhead.
Full-inference zkVM [2]. A SNARK circuit over the complete transformer forward pass provides the strongest software G2 guarantee—computational integrity without TEE hardware—by cryptographically binding the agent’s execution to its registered code commitment $C_{P}=\mathrm{CommitProg}(P)$. This path is appropriate for high-assurance scenarios (e.g., financial settlement, regulated environments) where the proof generation overhead (${\sim}120$ s per interaction) is acceptable and provider re-execution cooperation is unavailable. Unlike the software replay path, this path does not require model re-execution by the verifier; the proof is self-contained.
All four paths satisfy the same G2 governance contract. The framework is agnostic to which path is used; the choice is declared in the agent’s certificate ($\rho_{v}$) and recorded in the ledger, making the G2 implementation transparent to auditors and downstream agents.
4.2 Reproducibility as a Necessary Condition for Software G2
For the software path, a key observation applies: behavioral verifiability is impossible without reproducibility. If an agent’s outputs cannot be independently regenerated—even approximately—then no post-hoc mechanism can distinguish authorized behavior from unauthorized behavior. This is a fundamental constraint on any software-based verification scheme.
This observation has a prescriptive implication for secure agent system design: high-assurance deployments without TEE must be engineered for reproducible output. Concretely, this means using deterministic sampling (temperature==0, fixed seeds), preferring structured outputs (code, JSON, SQL) over free-form prose where security-critical decisions are involved, and selecting inference providers that guarantee bitwise or near-bitwise reproducibility [6]. Agents that cannot commit to any reproducibility level receive a trust tier downgrade (§4.5), making the security cost of non-reproducibility explicit in the governance framework. More broadly, by making reproducibility a prerequisite for software G2 compliance, our framework transforms it from an implementation detail into a security-critical model performance metric—with implications for how models are specified, evaluated, and deployed (§4.9).